Cost of car depends on km driven. Write lineqr equation expressing dollars per month as function of km per month if given $158 month if drive 300 km and $230 a month if drive 1500 km. Predict cost for 1000 ad 2000 km driven a month.
Answer by checkley71(8403)

SERTING UP X & Y VALUES FOR EACH SITUATION WE HAVE:
(158,300) & (230,1500)
NOW WE FIND THE SLOPE:
(1500-300)/(230-158)=1200/72=16.67
(1000-300)/(X-158)=16.67
16.67(X-158)=700
16.67X-2633.86=700
16.67X=700+2633.86
16.67X=3333.86
X=3333.86/16.67
X=$200 COST FOR DRIVING 1000KMs.
----------------------------------
(2000-300)/(X-158)=16.67
16.67(X-158)=1700
16.67X-2633.86=1700
16.67X=1700+2633.86
16.67X=4333.86
X=4333.86/16.67
X=$260 COST FOR DRIVING 2000KMs.